# BootOOD: Self-Supervised Out-of-Distribution Detection via Synthetic Sample Exposure under Neural Collapse

### v1

@@ 1 Introduction
Deep neural networks achieve impressive accuracy on large-scale image classification benchmarks, yet their predictions can be brittle when inputs deviate from the training distribution. In safety-sensitive applications such as autonomous driving, medical diagnosis, and open-world recognition, this vulnerability can lead to high-confidence errors that undermine reliability. Out-of-distribution (OOD) detection addresses this issue by equipping a classifier with the ability to detect inputs that differ from in-distribution (ID) training data and abstain or defer decisions when necessary [Hendrycks2017MSP, Guo2017Calibration, Hendrycks21Faces]. Despite intensive progress, reliably detecting OOD samples remains challenging, especially when OOD data are semantically similar to ID classes.
Early work on OOD detection focused on post-hoc scoring rules applied to a frozen classifier. The simplest baseline uses the maximum softmax probability (MSP) as an OOD score [Hendrycks2017MSP], and subsequent methods refine this idea through calibration [Guo2017Calibration], feature-space distances [Ren2018Mahalanobis], energy-based scores [Liu20Energy], gradient statistics [Huang2021GradNorm], activation rectification [Sun2021ReAct], nearest neighbors [Sun2022DeepKNN], sparsification [Sun2022DICE], or representation manipulation [Wang2022ViM, Song2022RankFeat, Djurisic2023ASH, Zhang2023SHE]. These post-hoc methods are attractive because they can be deployed on off-the-shelf classifiers without modifying training. However, large-scale evaluations have shown that they often saturate under realistic and fine-grained distribution shifts [Hendrycks2022ScalingOOD, Yang22OpenOODv1, Zhang24OpenOODv15, Bitterwolf2023NINCO], particularly when OOD samples share visual semantics with ID classes rather than being trivially far from the training distribution.
@@
We introduce BootOOD, a fully self-supervised OOD detection framework that is designed specifically for semantically similar OOD scenarios. The key idea is to relieve the main classifier from the burden of accommodating OOD data during training and to move OOD modeling into a lightweight auxiliary head that operates solely on feature norms. BootOOD bootstraps synthetic pseudo-OOD features by interpolating and perturbing ID representations that lie between collapsed class means, then trains an auxiliary radius-based classifier to regard these pseudo-OOD features as having shorter norms than ID features. This norm-based separation is a relaxed requirement compared to enforcing geometric orthogonality between ID and OOD, and it leverages the consistent norms induced by NC to define a simple radial decision rule.
@@ 2.4 Neural Collapse and Feature Norm for OOD Detection
Neural Collapse has emerged as a unifying perspective on the geometry of deep classifiers at convergence [Papyan2020NC]. Beyond explaining training dynamics, NC has motivated OOD and open-set recognition methods that exploit the structure of collapsed features and classifier weights. NECO [Ammar2024NECO], feature separation approaches [Wu2025SepLoss], and NC-inspired detectors [Liu24NC, Haas23LinkingL2NC] enforce geometric constraints on synthetic or real OOD samples, often by encouraging them to lie in subspaces orthogonal to ID class means or by manipulating feature norms. Park et al. [Park23FeatureNorm] further analyze the role of feature norms for OOD detection and show that norm statistics carry useful information about distributional shifts. These methods highlight that neural-collapse–induced structure can serve as a powerful inductive bias for OOD detection, yet many existing approaches still couple OOD constraints directly to the classifier and rely on strong angular-separation assumptions between ID and OOD samples [Ammar2024NECO, Wu2025SepLoss].
@@ 3.2 Pseudo-OOD Feature Generation
Given an ID minibatch with features $\{h_{i}\}_{i=1}^{B}$, we generate pseudo-OOD features directly in the feature space. Two distinct features $(h_{i},h_{j})$ are randomly paired, and we sample $\lambda\sim\mathrm{Beta}(\alpha,\alpha)$ to form a mixed representation
$h_{\mathrm{ood}}=\lambda h_{i}+(1-\lambda)h_{j},\qquad i\neq j.$ (1)
@@ 25  Update $\theta$, $\phi$, and $W$ using $\nabla\mathcal{L}$;
Why Mixup-Generated Pseudo-OOD Targets Near-OOD.
Near-OOD samples are known to lie in the low-density, between-class region of the ID feature space and often behave as interpolations between ID semantic manifolds [Sun2022DeepKNN, Park23FeatureNorm]. This region corresponds to the complement of high-density class neighborhoods within the convex hull:
$\mathrm{Conv}\big(\{z_{c}\}\big)\setminus\bigcup_{c}\mathcal{N}(z_{c}).$
Mixup as surrogate near-OOD. Feature-level mixup,
$z_{\mathrm{mix}}=\lambda z_{i}+(1-\lambda)z_{j},\qquad\lambda\sim\mathrm{Beta}(\alpha,\alpha),$
guarantees that $z_{\mathrm{mix}}\in\mathrm{Conv}(\{z_{c}\})$, and therefore places pseudo-OOD exactly in this between-class low-density region.
NC geometry. Under Neural Collapse, ID class means form a Simplex ETF [Papyan2020NC]; convex combinations lie on its faces, which are the natural decision boundaries. Thus mixup features fall near ETF faces—the same geometric zone where near-OOD typically appears.
Inference and OOD scoring.
@@ 4.1 Datasets, Protocols, and Metrics
For CIFAR-10 as ID, the near-OOD pool contains CIFAR-100 and Tiny-ImageNet (TIN) with overlapping classes removed [Krizhevsky2009CIFAR]; the far-OOD pool contains MNIST [deng2012mnist], SVHN [Netzer2011SVHN], Textures [Cimpoi2014DTD], and Places365 [Zhou2018Places]. For CIFAR-100 as ID, the near-OOD pool is CIFAR-10 and TIN, and the far-OOD pool matches the CIFAR-10 setting. For ImageNet-200 as ID, following OpenOOD, the near-OOD pool uses SSB-Hard and NINCO [Bitterwolf2023NINCO]; the far-OOD pool consists of iNaturalist [VanHorn2018iNat], Textures, and OpenImage-O [Wang2022ViM].
@@
We validate using multiple post-processors—entropy[Liu2023GEN], norm[Park23FeatureNorm, Yu2023FeatureNormBlock], ReAct[Sun2021ReAct], EBO[Liu20Energy], and MSP[Hendrycks2017MSP]. Because different datasets favor different scoring rules, we select the post-processor that achieves the best performance on the ID-only validation split, and use that fixed scorer for final testing. The backbone is unchanged at test time; only the scoring rule differs.
@@
To ensure Phase 1 reaches the Neural Collapse regime before Phase 2, we require training ID-only data beyond zero training error : NC typically appears around $\sim 70$ epochs on CIFAR-10, $\sim 80$ on CIFAR-100, and $\sim 100$ on ImageNet-200.
@@ 4.2 Baselines
We follow the OpenOOD v1.5 taxonomy and compare it with representative methods with respect to performance and recency of publication from three categories: post-hoc inference, outlier-free training, and outlier-exposed training. The complete list of methods and their respective settings is presented in the Appendix.
@@ 4.3 Main Results
We compare BootOOD against representative methods from the three standard OOD categories in OpenOOD: post-hoc methods, training-time methods without real OOD, and training-time methods with real OOD (four baselines per category). Following the OpenOOD protocol and its leaderboard focus, we report results on the near-OOD setting only, which is widely regarded as the most challenging regime. For CIFAR-10/100 and ImageNet-200, all numbers in Table 2 are averaged over 3 random seeds.
@@
Discussion of trends and failure modes.
Across datasets, the largest relative improvements occur on near-OOD, corroborating recent findings that near-OOD is intrinsically more challenging than far-OOD [Vaze2022OSR, Bitterwolf2023NINCO]. Our radius-based regularization most benefits categories with heavy inter-class confusion, where tightening ID manifolds reduces spurious high confidence on semantically similar OODs. Typical failure cases arise when near-OOD samples lie extremely close to decision boundaries (e.g., fine-grained bird species vs. ImageNet-200 classes); in those cases the score distributions can partially overlap, and improvements rely on careful tuning of the ramp schedule and margin. On far-OOD, improvements are consistent but smaller because separability is already easier.
@@
To inspect what BootOOD learns, we plot the distributions of (i) feature radius $\|\mathbf{z}\|_{2}$ and (ii) the maximum cosine to class weights $\max_{c}\cos(\mathbf{z},\mathbf{w}_{c})$ over ID test images and pseudo-OOD samples gathered at inference. As shown in Fig. 2 (left), ID features concentrate at larger radii while pseudo-OOD concentrate at smaller radii, with only a narrow overlap near the boundary; this matches the design of our radius self-supervision that encourages compact, larger-norm ID manifolds and discourages large norms for OOD-like points. In Fig. 2 (right), ID samples exhibit consistently higher maximal cosine to some class prototype than pseudo-OOD, indicating that ID features align more strongly with class directions whereas pseudo-OOD remain less aligned. Together, the two diagnostics corroborate the mechanism observed in the ablations: the radius terms provide the primary separability (clear norm gap), and the feature separation term further reduces alignment of pseudo-OOD with any class, which helps especially on near-OOD. We found the same qualitative pattern across seeds, and the operating thresholds selected on validation are therefore stable.
Figure 2: CIFAR-100 feature diagnostics: (left) radius $\|\mathbf{z}\|_{2}$; (right) max cosine to class weights $\max_{c}\cos(\mathbf{z},\mathbf{w}_{c})$. ID (blue) vs. pseudo-OOD (orange).
On the seemingly small cosine for ID.
Note that the ID histogram in Fig. 2(right) peaks around 0.30.40.3\!-\!0.4, which may look modest despite strong AUROC. This is expected because the class logit decomposes as $\ell_{c}=\mathbf{w}_{c}^{\top}\mathbf{z}=\|\mathbf{w}_{c}\|\,\|\mathbf{z}\|\,\cos\theta_{c}$. Our method primarily increases separability through the radius term: ID features have larger $\|\mathbf{z}\|$ while pseudo-OOD have smaller $\|\mathbf{z}\|$ (Fig. 2, left), so the dot product (and related energy scores) remains well separated even when $\cos\theta$ is moderate. Moreover, in high dimensions the maximum cosine over $100$ classes need not be large to be discriminative, and we do not enforce a cosine (unit-norm) classifier; hence the model leverages norm $\times$ moderate alignment rather than extreme alignment. This explains why AUROC is high while the ID cosine values are not; the norm gap is the main driver, with cosine providing complementary margins, especially near class boundaries.

### v2

@@ 1 Introduction
Deep neural networks achieve impressive accuracy on large-scale image classification benchmarks, yet their predictions can be brittle when inputs deviate from the training distribution. In safety-sensitive applications such as autonomous driving, medical diagnosis, and open-world recognition, this vulnerability can lead to high-confidence errors that undermine reliability. Out-of-distribution (OOD) detection addresses this issue by equipping a classifier with the ability to detect inputs that differ from in-distribution (ID) training data and abstain or defer decisions when necessary [Hendrycks2017MSP, Guo2017Calibration, Hendrycks21Faces]. Despite intensive progress, reliably detecting OOD samples remains challenging, especially when OOD data are semantically similar to ID classes [Yang2021SCOOD, DSAAhmed].
Early work on OOD detection focused on post-hoc scoring rules applied to a frozen classifier. The simplest baseline uses the maximum softmax probability (MSP) as an OOD score [Hendrycks2017MSP], and subsequent methods refine this idea through calibration [Guo2017Calibration], feature-space distances [Ren2018Mahalanobis], energy-based scores [Liu20Energy], gradient statistics [Huang2021GradNorm, abbas2025MedianofGradients], activation rectification [Sun2021ReAct], nearest neighbors [Sun2022DeepKNN], sparsification [Sun2022DICE], or representation manipulation [Wang2022ViM, Song2022RankFeat, Djurisic2023ASH, Zhang2023SHE]. These post-hoc methods are attractive because they can be deployed on off-the-shelf classifiers without modifying training. However, large-scale evaluations have shown that they often saturate under realistic and fine-grained distribution shifts [Hendrycks2022ScalingOOD, Yang22OpenOODv1, Zhang24OpenOODv15, Bitterwolf2023NINCO], particularly when OOD samples share visual semantics with ID classes rather than being trivially far from the training distribution.
@@
We introduce BootOOD, a fully self-supervised OOD detection framework that is designed specifically for semantically similar OOD scenarios. The key idea is to relieve the main classifier from the burden of accommodating OOD data during training and to move OOD modeling into a lightweight auxiliary head that operates solely on feature norms. BootOOD bootstraps synthetic pseudo-OOD features by interpolating ID representations that lie between collapsed class means, then trains an auxiliary radius-based classifier to regard these pseudo-OOD features as having shorter norms than ID features. This norm-based separation is a relaxed requirement compared to enforcing geometric orthogonality between ID and OOD, and it leverages the consistent norms induced by NC to define a simple radial decision rule.
@@ 2.4 Neural Collapse and Feature Norm for OOD Detection
Neural Collapse (NC) has emerged as a unifying perspective on the geometry of deep classifiers in the terminal phase of training [Papyan2020NC, Vignesh2023NCreview]. It characterizes a coupled set of phenomena where within-class features collapse to class means, centered class means approach a simplex equiangular tight frame (simplex ETF), last-layer weights align with these means up to scaling (self-duality), and the resulting classifier behaves similarly to a nearest-class-center rule in feature space. An important consequence of this geometry is the strong consistency of feature norms across classes, yielding near-spherical and class-balanced ID clusters [Vignesh2023NCreview].
Beyond explaining training dynamics, NC has motivated OOD and open-set recognition methods that exploit the structure of collapsed features and classifier weights. NECO [Ammar2024NECO], feature separation approaches [Wu2025SepLoss], and NC-inspired detectors [Liu24NC, Haas23LinkingL2NC] enforce geometric constraints on synthetic or real OOD samples, often by encouraging them to lie in subspaces orthogonal to ID class means or by manipulating feature norms. Park et al. [Park23FeatureNorm] further analyze the role of feature norms for OOD detection and show that norm statistics carry useful information about distributional shifts.
Complementary to NC-inspired geometry, between-class interpolation techniques such as Between-Class learning [tokozume2018between] and mixup [zhang2018mixup] synthesize samples that lie between ID classes. Such interpolations are known to populate low-density, decision-boundary-adjacent regions in representation space and can serve as structured surrogates for ambiguous or hard examples.
These lines of work highlight that neural-collapse–induced structure can serve as a powerful inductive bias for OOD detection, yet many existing approaches still couple OOD constraints directly to the classifier and rely on strong angular-separation assumptions between ID and OOD samples [Ammar2024NECO, Wu2025SepLoss].
Our differences.
@@ 3.1 Problem Setup
3.2 Neural Collapse as a Geometric Prior for BootOOD
NC essentials and geometry.
We build on the empirical observation that modern deep classifiers often enter a Neural Collapse (NC) regime in the terminal phase of training. In this regime, penultimate features concentrate around class means, while centered class means form an approximate simplex ETF and last-layer weights align with them (self-duality) [Papyan2020NC, Vignesh2023NCreview]. Concretely, letting $\mu_{c}\in\mathbb{R}^{d}$ denote the mean feature of class $c$ and $\tilde{\mu}_{c}$ its centered version,
$\tilde{\mu}_{c}$ $=\mu_{c}-\frac{1}{C}\sum_{j=1}^{C}\mu_{j},$ (1)
$\|\tilde{\mu}_{c}\|$ $\approx m,$
$\langle\tilde{\mu}_{i},\tilde{\mu}_{j}\rangle$ $\approx m^{2}\cos\theta,\quad i\neq j,$
$\cos\theta$ $=-\frac{1}{C-1},$
$w_{c}$ $\approx\alpha\,\tilde{\mu}_{c}.$
where $C$ is the number of classes, $m>0$ and $\alpha>0$ are global scales, and $w_{c}$ is the $c$-th classifier weight. A practically useful consequence is that ID feature norms become highly consistent across classes, producing near-spherical ID clusters [Vignesh2023NCreview].
Phase-1 $\rightarrow$ Phase-2 warm start.
BootOOD explicitly leverages this NC geometry. In Phase-1, we train a standard classifier until it approaches the NC regime, and estimate a global ID center $\mu$ and a reference radius $r_{\text{ref}}$ using exponential moving averages:
$\mu$ $\approx\mathbb{E}[h],$ (2)
$r_{\text{ref}}$ $\approx\mathbb{E}\bigl[\|h-\mu\|\bigr],$
where $h$ denotes penultimate features of ID samples. This yields a stable backbone geometry on which we attach an OOD-specific auxiliary head.
Why NC helps OOD detection.
Under NC, ID data already forms tight, near-spherical clusters with consistent radii. BootOOD converts this geometry into an inductive bias for OOD detection: pseudo-OOD features are driven towards smaller radii (norm separation) and non-ID directions (angular separation), while preserving the NC structure of the backbone and classifier. This enlarged margin improves standard post-hoc scores (MSP, entropy, feature norm) and aligns with evidence that NC-induced geometry benefits OOD and open-set recognition [Papyan2020NC, Vignesh2023NCreview, Ammar2024NECO].
3.3 Radius head and directional separation.
In Phase-2, we keep optimizing the backbone and the classifier using the standard in-distribution objective, while attaching an auxiliary radius head $h_{\mathrm{rad}}:\mathbb{R}^{d}\to\mathbb{R}^{K}$ to model the radial structure of features relative to the global ID center $\mu$. Importantly, the classifier continues to be trained by the ID classification loss; OOD-specific objectives are designed such that they do not directly alter the decision head.
Given a feature $h$, we consider its radius $\|h-\mu\|$ and define target radii
$\{\rho_{k}\}_{k=1}^{K}$ $\subset[R_{\min},\,\rho_{\max}],$ (3)
$\rho_{\max}$ $=(1-\gamma)\,r_{\mathrm{ref}},$
linearly spaced between $R_{\min}$ and $\rho_{\max}<r_{\mathrm{ref}}$. Here $\gamma\in(0,1)$ is a margin fraction that enforces pseudo-OOD features to lie strictly inside the typical ID radius.
The radius head is trained using a combination of (i) a $K$-way radial classification loss and (ii) a regression term aligning $\|h-\mu\|$ to its designated target radius $\rho_{k}$, weighted by $\lambda_{\mathrm{cls}}$ and $\lambda_{\mathrm{mse}}$, respectively.
To complement radial separation, we additionally penalize the cosine similarity between pseudo-OOD features and classifier weights $\{w_{c}\}$. Under the NC3 alignment in Eq. (1), this discourages pseudo-OOD features from aligning with the ID simplex-ETF directions, providing an angular margin without modifying the backbone classifier.
Overall, the auxiliary radius head and the separation objective act as training-time geometric regularizers that reshape the representation space beyond the ID manifold. Importantly, they do not participate in inference. At test time, the learned classifier is kept unchanged, and OOD detection is performed using standard post-hoc scoring functions on the resulting feature geometry.
@@ 3.4 Pseudo-OOD Feature Generation
Given an ID minibatch with penultimate features $\{h_{i}\}_{i=1}^{B}$ and labels $\{y_{i}\}_{i=1}^{B}$, we generate pseudo-OOD features directly in the feature space. Two features $(h_{i},h_{j})$ from different ID classes are randomly paired, and we sample $\lambda\sim\mathrm{Beta}(\alpha,\alpha)$ to form a mixed representation
$h_{\mathrm{ood}}=\lambda h_{i}+(1-\lambda)h_{j},\qquad y_{i}\neq y_{j}.$ (4)
@@
Why Mixup-Generated Pseudo-OOD Targets Near-OOD.
Near-OOD samples are known to lie in the low-density, between-class region of the ID feature space and often behave as semantic interpolations between ID manifolds [zhang2018mixup, tokozume2018between]. This region can be viewed as the complement of high-density class neighborhoods within the convex hull:
$\mathrm{Conv}\big(\{h_{c}\}\big)\setminus\bigcup_{c}\mathcal{N}(h_{c}).$
Feature-level mixup in Eq. (4) guarantees that $h_{\mathrm{ood}}\in\mathrm{Conv}(\{h_{c}\})$, placing pseudo-OOD features exactly in this between-class zone.
Connections to Between-Class learning and standard mixup.
Between-Class (BC) learning [tokozume2018between] mixes samples from different ID classes and regresses the mixing ratio, encouraging representations to populate structured division points between classes. Standard mixup [zhang2018mixup] performs convex combinations (typically in input space) and uses soft labels for vicinal risk minimization. In contrast, BootOOD applies a mixup-style operator directly in feature space and never treats mixed features as ID: we do not interpolate labels, and all mixed features are explicitly tagged as pseudo-OOD and supervised only by the auxiliary radius head. This preserves standard ERM on genuine ID samples while injecting structured near-OOD supervision without external OOD data.
NC geometry and norm shrinkage under simplex ETF.
Under Neural Collapse, centered class means form a simplex ETF with equal norm [Papyan2020NC, Vignesh2023NCreview]. Let $\mu_{c}=\mathbb{E}[h\,|\,y=c]$ and define the global mean $\mu=\frac{1}{C}\sum_{j=1}^{C}\mu_{j}$, with centered means $\tilde{\mu}_{c}=\mu_{c}-\mu$. In the NC regime, $\|\tilde{\mu}_{c}\|\approx m$ and $\langle\tilde{\mu}_{i},\tilde{\mu}_{j}\rangle\approx m^{2}\cos\theta$ for $i\neq j$, with $\cos\theta=-\frac{1}{C-1}$.
Consider cross-class mixup at the (centered) class-mean level:
$\tilde{h}_{\mathrm{mix}}=\lambda\,\tilde{\mu}_{i}+(1-\lambda)\,\tilde{\mu}_{j},\qquad i\neq j,\ \lambda\in(0,1).$
Then
$\|\tilde{h}_{\mathrm{mix}}\|^{2}$ $=m^{2}\bigl[\lambda^{2}+(1-\lambda)^{2}+2\lambda(1-\lambda)\cos\theta\bigr]$ (5)
$=m^{2}\bigl[1-2\lambda(1-\lambda)(1-\cos\theta)\bigr].$
Because $1-\cos\theta>0$ for a simplex ETF, Eq. (5) implies $\|\tilde{h}_{\mathrm{mix}}\|<m$ whenever $\lambda\in(0,1)$. Hence, between-class mixup features lie on simplex faces and exhibit strictly smaller centered norms than class centers. Moreover, with $\lambda\sim\mathrm{Beta}(\alpha,\alpha)$, larger $\alpha$ concentrates $\lambda$ near $1/2$, pushing typical mixup features deeper into the between-class faces and further shrinking their centered norms.
From between-class geometry to radius supervision.
Feature norm is a class-agnostic confidence proxy for OOD detection [Park23FeatureNorm]. Combining this with the ETF analysis above, mixup-generated features are simultaneously (i) between-class in direction and (ii) smaller in radius relative to the global center $\mu$. Our auxiliary radius head exploits exactly these cues: it assigns mixed features to inner radius shells (smaller target radii) while keeping ID features near the reference radius estimated from the NC backbone. Compared to BC learning, which regresses the mixing ratio, we only require pseudo-OOD to occupy low-norm, between-class shells—a relaxed constraint that is particularly suitable for semantically similar near-OOD settings.
Connection to feature-norm OOD scoring.
Since feature norm correlates with confidence [Park23FeatureNorm], mixup-induced norm shrinkage naturally strengthens norm-based OOD scoring and complements MSP/Entropy. Empirically, this is consistent with NC-inspired post-hoc detectors such as NECO [Ammar2024NECO].
@@ 26  Update $\theta$, $\phi$, and $W$ using $\nabla\mathcal{L}$;
3.10 Inference and OOD Scoring
@@ 4.1 Datasets, Protocols, and Metrics
For CIFAR-10 as ID, the near-OOD pool contains CIFAR-100 and Tiny-ImageNet (TIN) with overlapping classes removed [Krizhevsky2009CIFAR, Yang2021SCOOD]; the far-OOD pool contains MNIST [deng2012mnist], SVHN [Netzer2011SVHN], Textures [Cimpoi2014DTD], and Places365 [Zhou2018Places]. For CIFAR-100 as ID, the near-OOD pool is CIFAR-10 and TIN, and the far-OOD pool matches the CIFAR-10 setting. For ImageNet-200 as ID, following OpenOOD, the near-OOD pool uses SSB-Hard [Vaze2022OSR] and NINCO [Bitterwolf2023NINCO]; the far-OOD pool consists of iNaturalist [VanHorn2018iNat], Textures, and OpenImage-O [Wang2022ViM].
@@
We validate using multiple post-processors—entropy[Liu2023GEN], norm[Park23FeatureNorm, Yu2023FeatureNormBlock], ReAct[Sun2021ReAct], EBO[Liu20Energy], and MSP[Hendrycks2017MSP]. Because different datasets favor different scoring rules, we select the post-processor that achieves the best performance on the ID-only validation split, and use that fixed scorer for final testing, as shown in Table 1. The backbone is unchanged at test time; only the scoring rule differs. This protocol avoids tuning to a single scoring rule and provides a fair comparison across datasets.
@@
Our Phase-I training is designed to enter the terminal phase of training (TPT), where Neural Collapse (NC) is known to reliably emerge [Papyan2020NC, Vignesh2023NCreview]. To this end, we continue ID-only training beyond zero classification error before activating Phase-II regularization, ensuring that the feature geometry has reached a stable NC regime. Empirically, NC typically appears around $\sim 70$ epochs on CIFAR-10, $\sim 80$ epochs on CIFAR-100, and $\sim 100$ epochs on ImageNet-200.
@@ 4.2 Baselines
We follow the OpenOOD v1.5 taxonomy and compare it with representative methods with respect to performance and recency of publication from three categories: post-hoc inference, outlier-free training, and outlier-exposed training. The complete list of methods is presented in the Appendix.
@@ 4.3 Main Results
We compare BootOOD against representative methods from the three standard OOD categories in OpenOOD: post-hoc methods, training-time methods without real OOD, and training-time methods with real OOD (four baselines per category). Following the OpenOOD protocol and its leaderboard focus, we report results on the near-OOD setting only, which is widely regarded as the most challenging regime [Yang22OpenOODv1, Zhang24OpenOODv15, Yang2021SCOOD, DSAAhmed]. For CIFAR-10/100 and ImageNet-200, all numbers in Table 2 are averaged over 3 random seeds.
@@
Discussion of trends.
Across datasets, the largest relative improvements consistently occur on near-OOD, corroborating prior observations that near-OOD detection is intrinsically more challenging than far-OOD [Yang22OpenOODv1, Zhang24OpenOODv15, Yang2021SCOOD, DSAAhmed]. Our radius-based regularization particularly benefits settings with strong inter-class semantic similarity, where tightening ID feature manifolds reduces spurious over-confidence on near-OOD samples. On far-OOD, improvements are more modest but stable, as the underlying separability is already high for most methods.
@@
To inspect what BootOOD learns, we analyze the distributions of (i) feature radius $\|\mathbf{z}\|_{2}$ and (ii) the maximum cosine to class weights $\max_{c}\cos(\mathbf{z},\mathbf{w}_{c})$, computed during the standard OpenOOD evaluation stage. The statistics are collected over ID test images and real OOD test images following the OpenOOD protocol. As shown in Fig. 2 (left), ID features concentrate at larger radii while OOD features concentrate at smaller radii, with only a narrow overlap; this matches the goal of our radius self-supervision, which encourages compact, larger-norm ID representations and suppresses large norms for OOD inputs. In Fig. 2 (right), ID samples exhibit consistently higher maximal cosine to some class prototype than OOD samples, indicating that ID features are more strongly aligned with class directions whereas OOD features remain less aligned. Together, the two diagnostics corroborate the mechanism observed in the ablations: the radius terms provide the primary separability via a clear norm gap, while the feature separation term further reduces spurious alignment of OOD samples with any class direction, which is particularly beneficial for near-OOD. We observe the same qualitative pattern across seeds, indicating that the geometry induced by BootOOD is stable under the evaluation protocol.
Figure 2: CIFAR-100 feature diagnostics: (left) radius $\|\mathbf{z}\|_{2}$; (right) max cosine to class weights $\max_{c}\cos(\mathbf{z},\mathbf{w}_{c})$. ID (blue) vs. OOD (orange).